You roll a 6-sided die.



What is P(not less than 2)?
Write your answer as a fraction or whole number.

Answer by math_helper(2461)

P(less than 2) = P('1' is rolled) = 1/6


P(not less than 2) = 1-P(less than 2) = 1-(1/6) = 5/6